# How Much Content Do LLMs Generate That Induces Cognitive Bias in Users?

### v1

@@ 5.4 Mitigation for Hallucination
Method $\varphi_{\mathrm{frame}}$downarrow $\bar{s}_{b}\uparrow$ $\bar{s}_{m}\uparrow$ $\bar{s}_{e}\uparrow$ $\psi_{\mathrm{pri}}\downarrow$ $\varphi_{\mathrm{frame}}$downarrow $\bar{s}_{b}\uparrow$ $\bar{s}_{m}\uparrow$ $\bar{s}_{e}\uparrow$ $\psi_{\mathrm{pri}}\downarrow$
@@
Table 1: Models introduce framing changes ($\varphi_{\mathrm{frame}}$) and filter/feature content based on position ($\psi_{\mathrm{pri}}$). $\varphi_{\mathrm{frame}}$ is the proportion of examples where the framing changed from the original to the summary (lower is better). $\bar{s}_{b}$, $\bar{s}_{m}$, and $\bar{s}_{e}$ represent the average cosine similarity between the summary and the beginning, middle, and end thirds of the source text, respectively. Higher values of $\bar{s}_{b}$, $\bar{s}_{m}$, and $\bar{s}_{e}$ indicate stronger content preservation from those respective sections. $\psi_{\mathrm{pri}}$ represents the percentage of examples with primacy bias as we defined in Equation 2.
6 Experiments
6.1 Datasets
News Interviews
News summaries should have broad coverage of the content, independent of content position, as all of what has been said is potentially relevant for a summary. As these interviews are supposedly an objective depiction of current states and affairs, a summary should not be more positive or negative than the original content. We use MediaSum, a media interview dataset for dialogue summarization from NPR and CNN Zhu et al. (2021) to measure framing and primacy effects. We randomly select 1000 data points with tokens $\leq 4000$ to have the full content fit into a model’s context.
Product Reviews
Previous work suggests that changing the framing of the product influences consumer purchase decisions Wei et al. (2024). To explore this effect in the context of LLMs outputs, we use the Amazon Reviews dataset Ni et al. (2019). It contains customer reviews from Amazon across various product categories, providing information on product quality, usability, and satisfaction Ni et al. (2019). We sample 1,000 random examples from the Electronics category, filtering for token lengths $\leq 4000$ to maintain full-context input for the model.
News Hallucination
We provide a dataset to quantify how models can discern factual information across different temporal boundaries and in the true and negated versions of a news piece.
Pre-Knowledge Cutoff Data. For the pre-knowledge cutoff evaluation, we leverage the “All the news” data all (2020), which includes articles collected from 27 major American publications from 2016 to 2020. This data is available for non-commercial research purposes. We sampled 2,700 random data points in our experiment.
Post-Knowledge Cutoff Data. To evaluate model performance on recent events (i.e., events not included in a model’s training data), we introduce a self-updating dataset, the NewsLensSync Akshaya (2025), specifically designed to measure ongoing hallucinations for content beyond models’ knowledge cutoff dates. We evaluate 2801 real news descriptions and their negated versions.
For both post- and pre-cutoff examples, we include the true news events and carefully crafted counterfactual versions. We used a transformer-based model to generate the counterfactual examples Anschütz et al. (2023).
6.2 Models
We use four open- and closed-source language models. Details on the models and all of our hyperparameter settings can be found in Appendix B.
@@ 7 Results
7.1 LLMs Alter the Framing of Reviews when Summarizing
@@
We observe that models change the text framing of their summary compared to the context in both media interviews and the Amazon dataset (Table 1). Most models (Llama and GPT-3.5-turbo) show a framing change fraction between 14.5% and 16%, whereas Phi-3-mini-4k-Instruct shows 34.5% framing changes. In Table 3, we show examples of a framing change between a summary and the ground-truth context. Table  2 includes a detailed evaluation of framing alterations (e.g., from neutral to positive, or positive to negative). In the MediaSum news interviews, most framing shifts are from neutral to negative between 7.3% and 8.6% and negative to neutral between 8.9% and 13.1%. In Amazon reviews, the most common shift is from positive to neutral between 6.2% and 7.4% for Llama and GPT-3.5-turbo models and 21.2% for Phi-3-mini-4k-Instruct. This suggests that the models tend to downplay positive sentiment in product reviews. This pattern is consistent across all evaluated models.
Original Review GPT-3.5-turbo Summary
@@
Table 3: The framing of this example was classified as neutral before the summarization task to negative after.
7.2 LLMs Summarization have Imbalanced Context Coverage
We find that on average, summaries exhibit greater alignment with the beginning of the source text compared to the middle and end sections (Table 1). For example, across all models and datasets, the average similarity to the beginning section ranges from 0.828 to 0.860, which is consistently higher than the similarity to the middle 0.822-0.842 and end 0.820-0.840 sections. This coverage imbalance is further quantified using our primacy bias score, which captures the proportion of examples where beginning content is prioritized by at least 5% over other middle section. For the Amazon Reviews dataset, $\psi_{\mathrm{pri}}$ ranges from 5.4% to 7.6%. Similarly, MediaSum summaries exhibit $\psi_{\mathrm{pri}}$ scores between 4.0% and 6.1%, which shows primacy bias even in a more structured news content.
7.3 LLMs Hallucinate Post Knowledge-Cutoff
Baseline results in Table 4 show consistent drops in strict accuracy $(s_{H}^{T})$ from pre-cutoff to post-cutoff data across all models. Llama-3-8B-Instruct declined from 26% to 21% ($\delta_{H}=5\%$), Llama-3.2-3B-Instruct dropped from 19% to 13% ($\delta_{H}=6\%$), and Phi-3-mini-4k-Instruct showed a $\delta_{H}=4\%$ decline.
Models frequently hallucinate or misclassify new events, highlighting their inability to distinguish true from false statements without recent training data. Llama models averaged a $\delta_{H}=5\%$ decline despite individual accuracies above 40% on both real and falsified news. Llama-3-8B-Instruct performed best on true events across both settings, while Phi-3-mini-4k-Instruct showed poor performance with only 8% strict accuracy $(s_{H}^{T})$ on real news data.
@@ 7.3 LLMs Hallucinate Post Knowledge-Cutoff
Table 4: Accuracies on real-news items and their falsified counterparts, evaluated before and after each model’s training cutoff date. Strict Accuracy $s_{H}^{T}(D_{j})$ counts only predictions where the model gets both members of each (True, False) pair correct.
@@ 7.4 Special Biases Require Special Mitigations
The 18 mitigations we study (definitions in Appendix C) do not alleviate all biases simultaneously. Each method involves trade-offs among all biases, and its effectiveness varies depending on the model and corpus. For example, techniques aimed at reducing position-based salience (e.g., primacy bias) often work by redistributing attention across the input. However, this can disrupt how the model maintains consistent sentiment in its responses.
Method $\bar{s}$ $\psi_{\mathrm{pri}}$$\varphi_{\mathrm{frame}}$§
@@
Table 5: Key mitigation results on Amazon Reviews for Llama-3-8B-Instruct. Mean content-coverage similarity $(\bar{s}_{b}+\bar{s}_{m}+\bar{s}_{e})/3$. Primacy-bias score. §Framing-change rate. Lower is better for $\psi_{\mathrm{pri}}$ and $\varphi_{\mathrm{frame}}$. Behavior on other corpora/models follows a similar qualitative pattern; see full results in Tab. 9.
7.4.1 Primacy Bias and Coverage
@@
The Weighted Summaries method, which assigns a fixed token budget to each text section, consistently increases overall content coverage. For instance, with Llama-3-8B-Instruct on Amazon Reviews, it boosts average coverage similarity from a baseline of 0.843 to 0.910. However, this comes at the cost of worsening the primacy bias score ($\psi_{\mathrm{pri}}$), which increases from 7.0% to 15.1%.
@@
Mirostat Decoding also lifts coverage; it is the only method that cuts $\psi_{\mathrm{pri}}$ on the small Phi-3 model (-1.2 pp) but worsens it on larger models.
@@
Position-Invariant Shuffle is useful diagnostically because it helps confirm that the model relies on word order. By shuffling sentences randomly, we remove positional cues; the resulting change in output shows the model’s sensitivity to input structure. While this method increases overall coverage (e.g., from 0.843 to 0.891 for Llama-3-8B), it degrades other metrics, increasing the primacy score $\psi_{\mathrm{pri}}$ to 11.5% and the framing change rate $\varphi_{\mathrm{frame}}$ to 21.2%.
7.4.2 Framing Consistency
@@
Weighted Token Decoding proves most effective for this bias. By down-weighting negative lexemes during decoding, it delivers the lowest framing-change fraction ($\varphi_{\mathrm{frame}}$) for Llama-3-8B, reducing it by 0.9 percentage points to 13.6%. However, this targeted intervention significantly worsens primacy bias, with $\psi_{\mathrm{pri}}$ increasing from 7.0% to 18.4%.
@@
Prompt-only nudges. Lightweight prompts such as Self-Awareness or Chain-of-Thought alter $\varphi_{\mathrm{frame}}$ by only a few percentage points in most settings. However, the Self-Awareness variant gives the best framing result on MediaSum for Llama-3-8B-Instruct, reducing $\varphi_{\mathrm{frame}}$ to 21.0% from a 21.9% baseline at a negligible computational cost.
7.4.3 Hallucination
Prompt Calibration using Chain-of-Thought. This mitigation strategy improved the strict accuracies for Llama-3-8B-Instruct by 1% on both pre and post cut-off data.
For Llama-3.2-3B-Instruct, the strict accuracy dropped by 2% on pre cutoff data and there was no significant improvement when applied on post cut-off data.
Phi-3-mini-4k-Instruct showed the most dramatic degradation. Real news accuracy plummeted from 14% to 3% on pre-cutoff data. Post-cutoff data showed similar degradation in real news detection while maintaining high falsified news detection.
The approach appears to work reasonably well for larger models like Llama-3-8B-Instruct but causes significant performance degradation in smaller models like Phi-3-mini-4k-Instruct, consistent with known limitations in factual recall for smaller language models Abdin et al. (2024).
@@ 7.4.3 Hallucination
Knowledge Boundary Awareness. Making Llama-3-8B-Instruct aware of its knowledge boundaries showed some significant improvements by 15% in falsified news detection and maintained it’s strict accuracy close to the baseline in pre-cutoff data. In case of Llama-3.2-3B-Instruct and Phi-3-mini-4k-Instruct models, the performance declined in pre- and post-cutoff date data.
@@
Epistemic Tagging. This method performed the best across all models. It associates confidence level during the fact checking mechanism (i.e., highly or less confident) alongside answering whether an event has occurred or not (Table 6)
@@
Phi-3-mini-4k-Instruct was able to outperform with a steep increase in the strict accuracy by 21%. The Llama-3.2-3B-Instruct also performed well with strict accuracies improved by 8% and a balanced independent accuracies for real and falsified news data above 50%. On the other hand, Llama-3-8B-Instruct’s accuracy in identifying real news improved and maintained the strict accuracy around the baseline.
@@ 8 Conclusion
When LLMs modify the framing or emphasis of the content they process, they can inadvertently shape human perception and decision-making. In this work, we quantify how much LLMs alter original content in a summarization and news fact-checking task. We find that models, on average, introduce framing bias to the user in 21.86% of instances, measured by sentiment shifts between the source text and the generated summary and primacy bias in 5.94% of the cases, where summaries disproportionately reflect the beginning of the original text. The models also, on average, hallucinate on post-knowledge-cutoff data questions in 57.33% of instances.
Across the mitigation strategies evaluated, the effectiveness of each approach depends on the model type and the bias targeted. Weighted Summaries and Mirostat decoding showed promise in reducing framing changes and positional biases, particularly in smaller and mid-sized models. Epistemic Tagging consistently improved factual reliability on post-knowledge-cutoff knowledge.
This paper takes a step toward a careful analysis and mitigation of content alteration to avoid the risk of LLMs introducing systemic biases into decision-making processes across various domains, including media, education, and public policy.

### v2

@@ 6.4 Hallucination Mitigation
Method $\varphi_{\mathrm{frame}}$ $\downarrow$ $\bar{s}_{b}\uparrow$ $\bar{s}_{m}\uparrow$ $\bar{s}_{e}\uparrow$ $\psi_{\mathrm{pri}}\downarrow$ $\varphi_{\mathrm{frame}}$ $\downarrow$ $\bar{s}_{b}\uparrow$ $\bar{s}_{m}\uparrow$ $\bar{s}_{e}\uparrow$ $\psi_{\mathrm{pri}}\downarrow$
@@
Qwen3-4B-Instruct 27.8% 0.921 0.896 0.893 19.6% 45.9% 0.898 0.879 0.869 13.0%
Gemma-3-27B-IT 26.5% 0.924 0.896 0.892 23.7% 41.9% 0.902 0.879 0.868 17.6%
Table 1: Models introduce framing changes ($\varphi_{\mathrm{frame}}$) and filter content based on position ($\psi_{\mathrm{pri}}$). $\varphi_{\mathrm{frame}}$ is the proportion of examples where the framing changed from the original to the summary (lower is better). $\bar{s}_{b}$, $\bar{s}_{m}$, and $\bar{s}_{e}$ represent the average cosine similarity between the summary and the beginning, middle, and end thirds of the source text, respectively (higher is better). $\psi_{\mathrm{pri}}$ is the percentage of examples with primacy bias as defined in Equation 2 (lower is better). All differences between $\bar{s}_{b}$ and $\bar{s}_{m}$ are statistically significant ($p<.001$).
@@ 7 Results
7.1 LLMs Alter Framing During Summarization
@@
We observe that models alter the framing of summaries, compared to the source content, in both media interviews and the Amazon dataset (Table 1). GPT-3.5-turbo and Llama models exhibit a framing change fraction between 14.5% and 16%, whereas Gemma-3-2-27B-IT, Qwen3-4B-Instruct, and Phi-3-mini-4k-Instruct show framing changes between 26.5% and 45.9%. In Table 3, we show an example of a framing change between a summary and the ground-truth context. Table 2 provides a detailed evaluation of framing alterations (e.g., from neutral to positive, or positive to negative). In the MediaSum news interviews, most framing shifts occur between neutral and negative, ranging from 7.3% to 8.6% for GPT-3.5-turbo and Llama models, and reaching 30.8% to 31.4% for Qwen3-4B-Instruct and Gemma-3-27B-IT. Shifts from negative to neutral range from 2.9% to 13.1% across models. In Amazon reviews, the most common shift is from positive to neutral, ranging from 6.2% to 7.4% for the Llama and GPT-3.5-turbo models, and 21.2% for Phi-3-mini-4k-Instruct. Qwen3-4B-Instruct and Gemma-3-27B-IT show higher rates of neutral-to-positive shifts (14.4% to 15.3%) compared to other models, suggesting different framing patterns across model families.
Original Review Summary
@@
Table 3: Example of framing shift: from neutral (original text) to negative (summary by GPT-3.5-turbo).
7.2 LLM Summaries have Imbalanced Context Coverage
We find that, on average, summaries align more closely with the beginning of the source text than with the middle and end sections (Table 1). For example, across all models and datasets, the average similarity to the beginning segment is consistently higher than the similarity to the middle and end segments. Paired t-tests between the beginning-summary similarity ($s_{b}$) and middle-summary similarity ($s_{m}$) are statistically significant across all evaluated models and datasets ($p<.001$). The detailed test statistics are provided in Appendix A.5. We further quantify this coverage imbalance using our primacy bias score. For the Amazon Reviews dataset, $\psi_{\mathrm{pri}}$ ranges from 5.4% to 23.7%. Similarly, MediaSum summaries exhibit $\psi_{\mathrm{pri}}$ scores between 4.0% and 17.6%, which shows primacy bias in structured news content.
7.3 LLMs Hallucinate on Post-Knowledge-Cutoff Data
Table 4 shows a consistent decrease in strict accuracy $(s_{H}^{T})$ from pre-cutoff to post-cutoff data across all models. Llama-3-8B-Instruct declined from 26% to 21% ($\delta_{H}=5\%$), Llama-3.2-3B-Instruct dropped from 19% to 13% ($\delta_{H}=6\%$), and Phi-3-mini-4k-Instruct showed a $\delta_{H}=4\%$ decline.
The consistently low strict accuracy, compared to actual news and falsified news accuracy, highlights a critical limitation: the persistent inability to reliably differentiate fact from fabrication. Strong performance on isolated real or falsified news does not guarantee that models can reliably distinguish truth from falsehood.
@@ 7.3 LLMs Hallucinate on Post-Knowledge-Cutoff Data
Table 4: Accuracies on real news items and their falsified counterparts, evaluated before and after each model’s training cutoff date. Strict Accuracy $s_{H}^{T}(D_{j})$ counts predictions where the model gets both members of each (True, False) pair correct.
@@ 7.4 Special Biases Require Special Mitigations
We discuss trade-offs among mitigation methods for different biases, and how their effectiveness varies depending on the model and corpus. For example, techniques aimed at reducing position-based salience (e.g., primacy bias) often work by redistributing attention across the input. However, this can disrupt the model’s ability to maintain consistent sentiment in its responses. All definitions of the 18 mitigation approaches we tested are in Appendix A.6.
Method $\bar{s}$ $\psi_{\mathrm{pri}}$% $\varphi_{\mathrm{frame}}$% §
@@
Table 5: Key mitigation results on Amazon Reviews for Llama-3-8B-Instruct. Mean content-coverage similarity $(\bar{s}_{b}+\bar{s}_{m}+\bar{s}_{e})/3$. Primacy-bias score. §Framing-change rate. Lower is better for $\psi_{\mathrm{pri}}$ and $\varphi_{\mathrm{frame}}$. Behavior on other corpora/models follows a similar qualitative pattern; see full results in Table A6.
7.4.1 Primacy Effect and Coverage
@@
The Weighted Summaries method, which assigns a fixed token budget to each text segment, consistently increases overall content coverage. For instance, with Llama-3-8B-Instruct on Amazon Reviews (Table 5), it increases average coverage similarity from a baseline of 0.843 to 0.910. However, this comes at the cost of an increased primacy bias score ($\psi_{\mathrm{pri}}$), which increases from 7.0% to 15.1%.
@@
Mirostat Decoding also improves coverage (Table A6); it is the only method that reduces $\psi_{\mathrm{pri}}$ (primacy bias) on the smaller Phi-3 model by 1.2 percentage points (from 5.4% to 4.2%) but worsens it on larger models (e.g., from 7.0% to 15.5% on Llama-3-8B for Amazon).
@@
Position-Invariant Shuffle is useful diagnostically because it helps confirm that the model relies on word order. By randomly shuffling sentences, we remove positional cues; the resulting change in output reveals the model’s sensitivity to input structure. While this method increases overall coverage (e.g., from 0.843 to 0.891 for Llama-3-8B on Amazon Reviews; Table A6), it degrades other metrics, increasing the primacy score $\psi_{\mathrm{pri}}$ to 11.5% and the framing change rate $\varphi_{\mathrm{frame}}$ to 21.2%, and comes at the risk of losing temporal information.
7.4.2 Framing Effect and Consistency
@@
Weighted Token Decoding is the most effective method for mitigating framing bias. By downweighting negative lexemes during decoding, it achieves the lowest framing-change fraction ($\varphi_{\mathrm{frame}}$) for Llama-3-8B, reducing it by 0.9 percentage points to 13.6% (Table A6). However, this targeted intervention significantly worsens primacy bias, with $\psi_{\mathrm{pri}}$ increasing from 7.0% to 18.4%.
@@
Prompt-only approaches: Prompts such as Self-Awareness or Chain-of-Thought, alter $\varphi_{\mathrm{frame}}$ by only a few percentage points in most settings. However, the Self-Awareness variant yields the best framing result on MediaSum for Llama-3-8B-Instruct, reducing $\varphi_{\mathrm{frame}}$ to 21.0% from a 21.9% (Table A6) baseline at a negligible computational cost.
7.4.3 Prompt Calibration and Epistemic Tagging improve Hallucination
Applying Prompt Calibration with Chain-of-Thought enhances strict accuracy for Llama-3-8B Instruct but shows limited utility for others. It improves the detection of falsified news by 40% for both pre- and post-cut-off data.
@@ 7.4.3 Prompt Calibration and Epistemic Tagging improve Hallucination
Knowledge Boundary Awareness improves Llama-3-8B-Instruct’s performance by 8% and 21% in falsified news detection in pre and post cut-off data, respectively, alongside maintaining its strict accuracy close to the baseline. For the other models, performance declines, despite observing a significant improvement in the detection of falsified news data.
@@
Epistemic Tagging performs best across all models. It associates a confidence level during the fact-checking process (i.e., high or low) with its response about whether an event occurred (Table 6)
@@
We observe an increase in strict accuracy for Phi-3-mini-4k-Instruct by 21% on post cut-off data and by 19% on pre cut-off data. The Llama-3.2-3B-Instruct also shows improved accuracy of 8% on post-cut-off data and 7% on pre-cut-off data, with balanced independent accuracies above 50% for real and falsified news data.
@@
8 Impact on Human Decision-Making
When LLMs modify the framing or emphasize certain aspects of the content they process, they can inadvertently shape human perception and decision-making when humans are exposed to the content. In this section, we empirically demonstrate that human decisions change when LLMs process the original text and introduce undesired changes to it.
Experiment Design
Participants choose between two manufacturers of a product (e.g., a headlamp) based on product reviews. We probe 1000 reviews for items in the home, electronics, and kitchen categories from the Amazon-Reviews-2023 dataset Hou et al. (2024) and generate summaries using GPT-3.5. To analyze the worst-case scenario of how LLM content alteration can influence decision-making, we use GPT-4 to identify product types with the most extreme framing changes and manually select products to show to participants. Specifically, we select 10 product pairs where one product review shifts from negative/neutral to positive after summarization, while the other maintains neutral framing.
Participants are divided into two conditions and view 10 product pairs, either as original reviews or summaries (5 of each format per participant, alternating between conditions). We measure manufacturer selection rates based on the framing of reviews. Additionally, participants are asked to provide their Willingness to Pay Mitchell and Carson (2013) (WTP) for the product from the manufacturer they picked. WTP describes the maximum price someone would pay, knowing the cost of another product. For example, if the participant selects manufacturer $M$ whose product is described by a positive summary, they learn that the product from manufacturer $W$ costs $\mathdollar X_{W}$ and decide how much more over $\mathdollar X_{W}$ they would pay to $M$. This measure enables us to quantify the strength of the change in behavior induced by altering the content by the LLM.
Hypothesis
We hypothesize that participants will be more likely to choose a manufacturer when they read a positively framed LLM summary of their product compared to when they read the original review (which is neutral or negative). Additionally, we posit that their Willingness to Pay will be higher when reading the positively framed summaries.
Participants
We recruit 72 participants on Prolific (32-40 per condition, English-speaking, 33 men, 39 women, aged 19-75). Participants are presented with instructions about the study and take a short comprehension quiz at the beginning of the study, before proceeding to the decision-making stage. Based on this quiz, we accept 70 submissions (3% rejection rate). On average, the experiment takes 32 minutes and pays \$4.83, which includes \$3 base pay and a performance bonus for selecting the option that matches the majority opinion (encouraging participants to make decisions they believe reflect typical consumer preferences).
Figure 3: Mean selection rates for manufacturers depending on whether participants read the original review of their product (neutral or negative framing) or its summary (positive framing). Error bars display 95% confidence intervals and stars denote significance levels (* p < 0.05, ** p < 0.01, *** p < 0.001).
Results
We see that humans make different decisions, depending on the framing of the text they are exposed to. Choice rates differ significantly between framing conditions ($\chi^{2}$ = 78, df = 1, $p<0.001$). Fig.˜3 shows that participants picked manufacturers when their products were described by summaries (positive framing) in 83.7% of the cases compared to 52.3% of the cases if these products were described by original reviews (neutral or negative framing). The effect is robust across products, but differs in strength (Fig.˜3). Additionally, we find that reading the better-framed summaries increases participants’ willingness to pay by 4.5% with respect to the original product price ($t=4.406,p<0.001$ according to the linear mixed-effects model). Together, these results demonstrate that exposing participants to summarized content significantly affects whether they would purchase a product and what price they would pay for it. This demonstrates that when LLM summarization alters the text’s framing, it can have measurable consequences on human behavior.
@@ 9 Conclusion
This paper quantifies content changes in summarization and news fact-checking tasks performed by LLMs. Averaged across all tested models, models introduce framing bias to the user in 26.42% of instances, and show primacy bias in 10.12% of the cases, where summaries disproportionately reflect content from the beginning of the original text. Users are exposed to hallucinations on post-knowledge-cutoff content in 60.33% of instances ( averaged across all tested models). Through our user study, we find that model bias (specifically, framing bias that occurs in review summarization) influences people’s decisions. We observe that a positive framing bias increases the likelihood that people will purchase a given product and enhances their perception of its value.
By evaluating 18 mitigation strategies, we find that each approach’s effectiveness depends on the model and the targeted bias. For example, Weighted Summaries and Mirostat Decoding show reduced framing changes and primacy biases, particularly in smaller and mid-sized models. Epistemic Tagging consistently improved factual reliability regarding post-knowledge-cutoff data for smaller models.
Our paper represents a step toward careful analysis and mitigation of content alteration introduced by LLMs to humans, and provides insight into its effects, aiming to reduce the risk of systemic bias in decision-making across media, education, and public policy.